a giraffe is six times heavier than a lion. a zebra is one-quarter as heavy as a lion. together they are 450kg. how much does each animal weigh?

Found 3 solutions by mananth, rapaljer, Alan3354:
Answer by mananth(16949)

a giraffe is six times heavier than a lion. a zebra is one-quarter as heavy as a lion. together they are 450kg. how much does each animal weigh?
..
let lion weigh x kg
giraffe= 6x
zebra = x/4
Total = 450.
..
x+x/4 +6x=450
29x/4 =450
x=1800/29
lion = 62.06 kgs
giraffe= 372.6 kgs
zebra = 15.5
CHECK
62.06 +372.6+15.5=450.6kgs

Answer by rapaljer(4671)

Let x = weight of the lion, since this is at the end of each phrase of the problem.

6x= weight of the giraffe
(1/4)x= weight of the zebra

x+6x + (1/4)x=450

Multiply both sides by 4:
4x+ 24x+ 1x= 1800
29x=1800
x=62.07 lbs approximately = LION
6x=372.41 lbs approximately = GIRAFFE
1/4 x = 15.52 lbs approximately = ZEBRA

Sum = 450 pounds


You really need to see my own website for all the FREE resources that I have there, including my OWN explanation of Word Problems. To see my website, click on my tutor name "Rapaljer" anywhere in algebra.com. On my information page, click on the link to my website. Look for the link on my Homepage "Basic, Intermediate, and College Algebra: One Step at a Time." Choose "Basic Algebra", then "Chapter 1" and look for Sections 1.09 and 1.10, the topics about Word Problems . This is my OWN non-traditional explanation, complete with examples, exercises, and ALL of the answers, that my own students found much easier to understand than the regular textbooks. See also my "Math in Living Color" pages that go with this section. In these pages, I explain the solutions to many of the exercises IN LIVING COLOR!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus

Answer by Alan3354(69443)

Is it "6 times heavier", or six times as heavy?
---------------------------
7x = weight of the giraffe
x/4 = weight of the zebra
x + 7x + x/4=450
Multiply thru by 4
4x+ 28x+ 1x= 1800
33x=1800
x =~ 54.54 lbs (weight of lion)
7x =~ 381.82 lbs (giraffe)
x/4 =~ 13.36 lbs (zebra)
Totals 450 pounds